# Yifan Wang1∗ Jianjun Zhou123∗ Haoyi Zhu1 Wenzheng Chang1 Yang Zhou1

### v1

@@ A.3 Discussion for Predicted Pose Distribution
A.4 More Experiments on Scalability and Convergence Speed
To evaluate the scalability and convergence speed of our approach, we conducted further experiments using three different model sizes. The results in Fig. 9, plotted as metric curves against training epochs, demonstrate that our method achieves consistent performance gains across all scales. Furthermore, our models converge significantly faster than the baseline. These findings underscore the effectiveness and importance of our permutation-equivariant design.
Figure 9: Comparison of the PI model (solid colored lines) against the non-PI baseline (dashed lines) across three model scales. The metric shown is the average of Accuracy and Completeness on the 7-Scenes [27] and NRGBD [1] datasets. The shaded area visually represents the performance improvement gained by the PI model.

### v2

@@ A.3 Discussion for Predicted Pose Distribution
A.4 Comparison with VGGT
This section details an experiment designed solely for a fair comparison against VGGT [34]. A direct comparison is challenging because training our model from scratch with only its core objectives (camera poses and local pointmaps) leads to suboptimal convergence, whereas VGGT’s design incorporates a multi-task learning setup.
For a fair comparison, and similar to VGGT, we also predict a global pointmap to serve as a regularizer. We temporarily adapt our training to mirror VGGT’s methodology. We introduce an auxiliary head to predict a global pointmap relative to a reference frame, using a loss analogous to Eq. 5. We directly use the scale factor from the alignment of local pointmaps. Please note that the reference view is incorporated as context via cross-attention, exclusively within the global pointmap head. This head only serves as a regularization term and our final model remains fully permutation-equivariant.
We train both our adapted model and VGGT under these identical, multi-task conditions: from scratch (except for DINOv2 encoders) on the same data, at a $224\times 224$ resolution for 80 epochs (800 steps/epoch). We use the same data as described in Section 3.4.
Table 9: Comparison with VGGT when trained from scratch.
Method ETH3D 7-Scenes NRGB
Acc. $\downarrow$ Comp. $\downarrow$ Acc. $\downarrow$ Comp. $\downarrow$ Acc. $\downarrow$ Comp. $\downarrow$
VGGT 0.563 0.449 0.057 0.046 0.060 0.042
$\pi^{3}$ (ours) 0.418 0.266 0.059 0.071 0.052 0.035
As shown in Table 9, $\pi^{3}$ outperforms VGGT on two of the three benchmarks. It is important to mention that for this VGGT baseline, we did not utilize its tracking branch, as the official implementation did not provide clear instructions or clean code for its usage.